Write a rule that represents the function?


Please provide a detailed explanation with the answer.
Thank you very much

Found 2 solutions by TimothyLamb, Edwin McCravy:
Answer by TimothyLamb(4379)

---
the given data points fall on the same line.
you can verify this fact by plugging the following list of the given data points:
0,-8, 1,-7, 2,-6, 3,-5, 4,-4
---
into this linear regression solver:
https://sooeet.com/math/linear-regression.php
---
linear regression model:
y-intercept: -8
slope: 1
---
answer:
a rule that represents the function, is the slope-intercept form of the linear equation:
y = x - 8
---
you can also verify this linear relationship by calculating the slope between each two data-points like so:
m = (y2 - y1)/(x2 - x1)
all such slopes are 1 because all points are on the same line.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by Edwin McCravy(20056)

(0,-8), (1,-7), (2,-6), (3,-5), (4,-4)
(0,-8),  we observe that (-8)-(0) = -8
(1,-7),  we observe that (-7)-(1) = -8
(2,-6),  we observe that (-6)-(2) = -8
(3,-5),  we observe that (-5)-(3) = -8
(4,-4),  we observe that (-4)-(4) = -8

We therefore conclude that if we had (x,y),
we would observe that 

y-x = -8

  y = x-8

That's the rule.

Edwin